# ArchRAG: Attributed Community-based Hierarchical Retrieval-Augmented Generation

### v1

@@ Abstract.
Retrieval-Augmented Generation (RAG) has proven effective in integrating external knowledge into large language models (LLMs) for question-answer (QA) tasks. The state-of-the-art RAG approaches often use the graph data as the external data since they capture the rich semantic information and link relationships between entities. However, existing graph-based RAG approaches cannot accurately identify the relevant information from the graph and also consume large numbers of tokens in the online retrieval process. To address these issues, we introduce a novel graph-based RAG approach, called Attributed Community-based Hierarchical RAG (ArchRAG), by augmenting the question using attributed communities, and also introducing a novel LLM-based hierarchical clustering method. To retrieve the most relevant information from the graph for the question, we build a novel hierarchical index structure for the attributed communities and develop an effective online retrieval method. Experimental results demonstrate that ArchRAG outperforms existing methods in terms of both accuracy and token cost.
Retrieval-Augmented Generation, Graph-based RAG
1. Introduction
Large Language Models (LLMs) like GPT-4  (Achiam et al., 2023), Qwen2.5  (Yang et al., 2024), and LLaMA3.1  (Dubey et al., 2024) have received tremendous attention from both industry and academia (Huang and Huang, 2024; Liu et al., 2024c; Wang et al., 2024b; Zheng et al., 2024; Li et al., 2023; Nie et al., 2024; Ghimire et al., 2024; Wang et al., 2024c). Despite their remarkable success in question-answering (QA) tasks, they may still generate wrong answers due to a lack of domain-specific, real-time updated, and proprietary knowledge outside their pre-training corpus (Peng et al., 2024). To enhance the trustworthiness and interpretability of LLMs, Retrieval-Augmented Generation (RAG) methods (Fan et al., 2024; Gao et al., 2023; Hu and Lu, 2024; Huang and Huang, 2024; Wu et al., 2024a; Yu et al., 2024; Zhao et al., 2024) have emerged as a core approach, which often retrieve relevant information from documents, relational data, and graph data to facilitate the QA tasks.
The state-of-the-art RAG approaches often use the graph data as the external data since they capture the rich semantic information and link relationships between entities. Given a question $Q$, the key idea of graph-based RAG is to retrieve relevant information (e.g., nodes, subgraphs, or textual information) from the graph, incorporate them with $Q$ as the prompt, and feed them into the LLM, as illustrated in Figure 1. As shown in the literature, a few recent graph-based RAG methods have emerged for both abstract question (Edge et al., 2024; Guo et al., 2024; Sarthi et al., 2024) and specific question (Guo et al., 2024; Sarthi et al., 2024; Wu et al., 2024b; Wang et al., 2024a; Li et al., 2024; Gutiérrez et al., 2024), where the former is more conceptual, encompassing broader topics, summaries, or overarching themes (e.g., “What are the potential impacts of LLMs on education?”), while the latter is detail-oriented and typically references specific entities within the graph (e.g., “Who has won the Turing Award in 2024?”) (Guo et al., 2024). Technically, the abstract questions need a comprehensive understanding to extract high-level insights, while the specific questions require multi-hop reasoning to connect discrete information. Table 1 summarizes the key characteristics of three representative graph-based RAG methods, i.e., GraphRAG (Edge et al., 2024), HippoRAG (Gutiérrez et al., 2024), and LightRAG (Guo et al., 2024).
@@ 1. Introduction
Figure 1. The general workflow of graph-based RAG, which retrieves relevant information (e.g., nodes, subgraphs, or textual information) to facilitate the LLM generation.
Specifically, Microsoft first proposes a graph-based RAG system called GraphRAG (Edge et al., 2024), which first builds a knowledge graph (KG) by extracting entities and relationships from the external corpus, then employs the Leiden method (Traag et al., 2019) to detect communities and generates a summary for each community using LLMs. When answering abstract questions, it uses a Global Search method, which traverses each community and uses LLMs to retrieve relevant communities to answer the questions; for specific questions, it presents a Local Search method, which retrieves entities, identifies the relevant chunk texts and low-level communities to augments the response. LightRAG (Guo et al., 2024) addresses abstract questions by adding keywords to relations in the constructed KG, and during querying, it uses LLMs to extract keywords and searches for the most relevant entities and relationships, which are then combined with the corresponding chunks for generation. HippoRAG (Gutiérrez et al., 2024) employs LLMs to extract tuples from the corpus to build a KG and connect edges between similar entities in KG to improve the quality of the KG. To answer specific questions requiring multi-hop reasoning, HippoRAG utilizes the Personalized PageRank (PPR) (Haveliwala, 2002) to identify the most relevant entities from the KG and augment the question with chunks containing these selected entities for LLM generation.
Table 1. Comparison of representative RAG methods and our ArchRAG.
RAG method Retrieval method Question Attributed community Hierarchical index Retrieval filtering
[…]
While these approaches have achieved good performance in many scenarios, they still have some following limitations: (1) GraphRAG detects communities by relying on the Leiden algorithm, which however only uses the graph structure without considering the rich semantic information carried by the nodes and edges. As a result, the detected communities often consist of various different themes, which leads to the poor quality of community summaries and further decreases its performance. Besides, it uses LLMs to analyze all the communities when answering abstract questions, which is both time- and tokens-consuming. For example, the Leiden algorithm can detect 2,984 communities in the Multihop-RAG dataset, and to answer 100 questions, Global Search costs approximately \$650 and 106 million tokens [^1] The cost of GPT-4o is \$10/M tokens for output and \$2.50/M tokens for input (for details, please refer to openai.com/api/pricing)., which is very token-consuming. (2) LightRAG still cannot accurately identify the important keywords from the questions, which greatly affects its performance, and it does not generate community summaries in advance, making it difficult for the LLM to generate comprehensive summaries of all retrieved information. (3) HippoRAG fails to capture relationships between different chunks, making it susceptible to the “lost in the middle” dilemma (Liu et al., 2024b), which significantly prevents its ability to answer abstract questions, as it struggles to synthesize information across multiple chunks or understand overarching themes.
To tackle the above limitations, in this paper we propose a novel graph-based RAG approach, called Attributed Community-based Hierarchical RAG (ArchRAG), by augmenting the question using attributed communities from the knowledge graph built on the external corpus. Unlike GraphRAG which detects communities purely using links, we propose to detect communities by exploiting both links and the attributes of nodes, and call them attributed communities. Intuitively, an attributed community represents a group of nodes that are not only densely connected but also share similar themes, which provides detailed information for answering specific questions, and its summary offers a condensed view (Sarthi et al., 2024; Angelidis and Lapata, 2018), enabling more focused engagement with the content and making it suitable for answering abstract questions. To effectively answer both abstract and specific questions, we then organize the attributed communities with summaries into a hierarchical structure, which provides different granularity of abstractions for the whole knowledge graphs. Besides, we develop a novel hierarchical data structure for searching the nearest neighbors and the corresponding efficient online retrieval method.
Specifically, to detect attributed communities that are organized in a hierarchical manner, we propose a novel LLM-based iterative framework, which can incorporate any existing attributed community detection method (Zhou et al., 2009; Traag et al., 2019; Von Luxburg, 2007; Xu et al., 2007; Grover and Leskovec, 2016). In each iteration, we detect the attributed communities by considering both links and node attributes using a community detection method and generate a summary for each attributed community by using LLMs. Afterwards, we build a new smaller attributed graph by abstracting each attributed community as a single node and then linking each pair of nodes together if their summary similarity exceeds a threshold. By iterating the above steps multiple times, we can obtain a set of attributed communities that can be organized into a hierarchical tree structure.
Given the attributed communities organized hierarchically, to answer a question, a naive approach is to enumerate all the attributed communities and find the most relevant ones. However, this approach, in line with that of GraphRAG, is very time and tokens-consuming. To alleviate these issues, we propose to build a novel hierarchical index structure by borrowing some ideas from HNSW (Malkov and Yashunin, 2018), which is a well-known approximate nearest neighbor (ANN) search method, and we call it C-HNSW (Community-based HNSW). C-HNSW index facilitates the fast retrieval of the $k$ most relevant elements from each layer of the hierarchy, optimizing the RAG process. Additionally, by using adaptive filtering, we identify the most relevant information, ensuring that complementary data are retrieved across all layers. Through hierarchical search, we achieve cost-efficient generation and optimize the retrieval process without sacrificing the quality of the generated output.
@@
We have extensively evaluated ArchRAG by conducting various experiments on real-world datasets, and the results show that ArchRAG outperforms existing approaches in answering both abstract and specific questions. Particularly, ArchRAG achieves a 10% high accuracy compared to state-of-the-art graph-based RAG methods when answering specific questions and also demonstrates significant improvement over others in answering abstract questions. Moreover, ArchRAG is very token-efficient, saving up to 250 times the token usage compared to GraphRAG (Edge et al., 2024).
@@
  • We present a novel graph-based RAG approach by using attributed communities that are organized hierarchically and detected by an LLM-based hierarchical clustering method.
@@
  • To index attributed communities, we propose a novel hierarchical index structure called C-HNSW and also develop an efficient online retrieval method.
@@
  • Extensive experiments show that ArchRAG is both highly effective and efficient, and achieves state-of-the-art performance in answering both abstract and specific QA tasks.
2. Preliminaries
@@
Figure 2. ArchRAG consists of two phases: offline indexing and online retrieval. For the online retrieval phase, we show an example of using ArchRAG to answer a question in the HotpotQA dataset.
Existing graph-based RAG methods often follow a general framework that leverages external structured knowledge graphs to improve contextual understanding of LLMs and generate more informed responses (Peng et al., 2024). Typically, it involves two phases:
  1. (1)
    Offline indexing: Building a KG $G(V,E)$ from a given corpus $D$ where each vertex represents an entity and each edge denotes the relationship between two entities, and constructing an index based on the KG.
  2. (2)
    Online retrieval: Retrieving the relevant information (e.g., nodes, subgraphs, or textual information) from KG using the index and providing the retrieved information to the LLM to improve the accuracy of the response.
3. Our Approach ArchRAG
Unlike GraphRAG, which purely uses links to generate communities, we propose to detect communities that can be organized into a hierarchical tree structure by exploiting both links and attributes in the knowledge graph, also called attributed communities. An attributed community is a group of entities with a summary, where the entities are not only densely connected but also share similar themes, providing detailed information for answering specific questions, and the summary can offer a condensed view for answering abstract questions since it enables more focused engagement with the entities’ attributes. Moreover, in the hierarchical tree structure, the lower-level communities and entities contain detailed knowledge from the KG, while higher-level communities provide global summaries, which naturally enable ArchRAG to address questions with different granularity in abstraction.
Moreover, to answer questions, we need to efficiently identify highly relevant attributed communities at different levels of the hierarchical structure. Considering that our hierarchical attributed communities and the HNSW index (Malkov and Yashunin, 2018) are similar in structure, we propose to map entities and attributed communities into high-dimensional nodes and then build a unified Community-based HNSW (C-HNSW) index for them, which allows us to efficiently identify the highly relevant attributed communities.
@@ 3. Our Approach ArchRAG
As illustrated in Figure 2, our proposed ArchRAG consists of two phases. In the offline indexing phase, ArchRAG first constructs a KG from the corpus, then detects attributed communities by a novel LLM-based hierarchical clustering method, and finally builds the C-HNSW index. During the online retrieval phase, ArchRAG first converts the question into a query vector, then retrieves relevant information from the C-HNSW index, and finally generates answers through adaptive filtering-based generation process which identifies the most relevant information.
3.1. Offline Indexing
3.1.1. KG construction.
@@ 3.1. Offline Indexing
Our first step is to build a KG by prompting the LLM to extract entities and relations from each chunk of the text corpus $D$. Specifically, all text contexts are segmented into chunks based on specified chunk length, enabling the LLM to extract entities and relations from each chunk using in-context learning (Brown, 2020), thus forming subgraphs. These subgraphs are then merged, with entities and relations that appear repeatedly across multiple subgraphs being consolidated by the LLM to generate a complete description. Finally, we get a KG, denoted by $G(V,E)$, where $V$ and $E$ are sets of vertices and edges, respectively, and each vertex and edge is associated with some textual attributes.
3.1.2. LLM-based hierarchical clustering.
@@
As aforementioned, a high-quality hierarchical attributed community structure provides different granularities of abstraction for the entire knowledge graph, which is beneficial for answering both abstract and specific questions. Therefore, we propose an iterative LLM-based hierarchical clustering framework that supports arbitrary graph augmentation (e.g., KNN connections and CODICIL (Ruan et al., 2013)) and graph clustering algorithms (e.g., weighted Leiden (Traag et al., 2019), weighted spectral clustering, and SCAN (Xu et al., 2007)). Specifically, we propose to augment the KG by linking entities if their attribute similarities are larger than a threshold (e.g., KNN connections), and then associate each pair of linked entities with a weight denoting their attribute similarity value. Next, we generate the attributed communities using any given graph clustering algorithm.
@@
Algorithm 1 shows the above iterative clustering process. Given a graph augmentation method and clustering algorithm, we perform the following steps in each iteration: (1) augmenting the graph (line 3); (2) computing the edge weights (lines 4-5); (3) clustering the augmented graph (line 6); (4) generating a summary for each community using LLM (line 7); and (5) building a new attributed graph where each node denotes an attributed community and two nodes are linked if their community members are connected (line 9). We repeat the iterations until the stopping condition (such as insufficient nodes or reaching the specified level limit) is met. Since each iteration corresponds to one layer, all the attributed communities $HC$ can be organized into a multi-layer hierarchical tree structure, denoted by $\Delta$, where each community in one layer includes multiple communities in the next layer.
Appendix A.1 provides a running example for Algorithm 1.
input : KG ${G}(V,E)$, augment function Aug, graph clustering algorithm GCluster, and termination condition $T$
@@ 1 $T\leftarrow$ False, ${HC}\leftarrow\emptyset$;
3       ${G}^{\prime}(V,E^{\prime})\leftarrow$ Aug(${G}(V,E)$);
@@
8      $C\leftarrow$ GCluster(${G}^{\prime}(V,E^{\prime})$);
9       for each $c\in C$ do  generate the summary of $c$ by LLM ;

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) has proven effective in integrating external knowledge into large language models (LLMs) for solving question-answer (QA) tasks. The state-of-the-art RAG approaches often use the graph data as the external data since they capture the rich semantic information and link relationships between entities. However, existing graph-based RAG approaches cannot accurately identify the relevant information from the graph and also consume large numbers of tokens in the online retrieval process. To address these issues, we introduce a novel graph-based RAG approach, called Attributed Community-based Hierarchical RAG (ArchRAG), by augmenting the question using attributed communities, and also introducing a novel LLM-based hierarchical clustering method. To retrieve the most relevant information from the graph for the question, we build a novel hierarchical index structure for the attributed communities and develop an effective online retrieval method. Experimental results demonstrate that ArchRAG outperforms existing methods in both accuracy and token cost.
Introduction
Retrieval-Augmented Generation (RAG) has emerged as a core approach for enhancing large language models (LLMs) by enabling access to domain-specific and real-time updated knowledge beyond their pre-training corpus (2024; 2023; 2024; 2024; 2024b; 2024; 2024). By improving the trustworthiness and interpretability of LLMs, RAG has been widely adopted across a broad range of applications (2024a; 2024a; 2024; 2024; 2023; 2024; 2024b). Current state-of-the-art RAG approaches often use graph-structured data as external knowledge, due to its ability to capture the rich semantics and relationships. Given a question $Q$, the key idea of graph-based RAG is to retrieve relevant information (e.g., nodes, subgraphs, or textual information) from the graph, incorporate them with $Q$ as the prompt, and feed them into the LLM, as illustrated in Figure 1. Several recent methods (2024; 2024; 2024; 2024a; 2024c; 2024; 2024) address two common question answering (QA) tasks: abstract questions, which require reasoning over high-level themes (e.g., “What are the potential impacts of LLMs on education?”), and specific questions, which focus on entity-centric factual details (e.g., “Who won the Turing Award in 2024?”).
@@ Introduction
Figure 1: The general workflow of graph-based RAG, which retrieves relevant information (e.g., nodes, subgraphs, or textual information) to facilitate the LLM generation.
In the past year, a surge of graph-based RAG methods (2025; 2024c; 2024; 2024; 2024; 2023) has emerged, each proposing different retrieval strategies to extract detailed information for response generation. Among them, GraphRAG (2024), proposed by Microsoft, is the most prominent and the first to leverage community summarization for abstract QA. It builds a knowledge graph (KG) from the external corpus, detects communities using Leiden (2019), and generates a summary for each community using LLMs. For abstract questions that require high-level information, it adopts a Global Search approach, traversing all communities and using LLMs to retrieve the most relevant summaries. In contrast, for specific questions, it employs a Local Search method to retrieve entities, relevant text chunks, and low-level communities, providing the multi-hop detailed information for accurate answers.
Although some methods claim that GraphRAG underperforms and is difficult to apply in practice (2024; 2024a), our re-examination shows that it is primarily constrained by the following limitations (L): L1. Low community quality: GraphRAG uses the Leiden (2019) algorithm to detect communities, but this approach relies solely on graph structure and ignores the rich semantics of nodes and edges. As a result, the detected communities often consist of different themes, which leads to the poor quality of community summaries and further decreases its performance. L2. Limited compatibility: While GraphRAG employs Global and Local Search strategies, each retrieves graph elements at only one granularity, making it inadequate for simultaneously addressing both abstract and specific questions and limiting its applicability in real-world open-ended scenarios. L3. High generation cost: Although GraphRAG performs well on abstract questions, analyzing all communities with LLMs is both time- and token-consuming. For example, GraphRAG detects 2,984 communities in the Multihop-RAG (2024) dataset, and answering just 100 questions incurs a cost of approximately \$650 and 106 million tokens [^1] The cost of GPT-4o is \$10/M tokens for output and \$2.50/M tokens for input (for details, please refer to OpenAI pricing)., which is an impractical overhead.
To tackle the above limitations of GraphRAG, in this paper, we propose a novel graph-based RAG approach, called Attributed Community-based Hierarchical RAG (ArchRAG). ArchRAG leverages attributed communities (ACs) and introduces an efficient hierarchical retrieval strategy to adaptively support both abstract and specific questions. To mitigate L1, we detect high-quality ACs by exploiting both links and the attributes of nodes, ensuring that each AC comprises nodes that are not only densely connected but also share similar semantic themes (2009). We further propose a novel LLM-based iterative framework for hierarchical AC detection, which can incorporate any existing community detection methods (2009; 2019; 2007; 2007; 2016). In each iteration, we detect ACs based on both attribute similarity and connectivity, summarize each AC using an LLM, and construct a higher-level graph by treating each AC as a node, connecting pairs with similar summaries. By iterating the above steps multiple times, we obtain a set of ACs that can be organized into a hierarchical tree structure.
To effectively address L2, we organize all ACs and entities into a hierarchical index and retrieve relevant elements from all levels to support both abstract and specific questions. Entities offer fine-grained details, while LLM-generated AC summaries capture relational structures and provide high-level condensed overviews (2024; 2018), making them suitable for both multi-hop reasoning and abstract insight extraction. To support efficient retrieval across levels, we propose C-HNSW (Community-based HNSW), a novel hierarchical index inspired by the HNSW algorithm (2018) for approximate nearest neighbor (ANN) search.
To mitigate the high generation cost caused by traversing all communities (L3), we propose a hierarchical search with adaptive filtering to efficiently select the most relevant ACs and entities while maintaining performance. Specifically, we design an efficient hierarchical retrieval algorithm over the proposed C-HNSW index, which supports top-$k$ nearest neighbor search across multiple levels, thereby facilitating access to multi-level relevant information. Furthermore, the adaptive filtering mechanism identifies the most informative results at each level, making the retrieved information complementary.
@@
We have extensively evaluated ArchRAG on real-world datasets, and the results show that it consistently outperforms existing methods in both abstract and specific QA tasks. ArchRAG achieves a 10% higher accuracy than state-of-the-art graph-based RAG methods on specific questions and shows notable gains on abstract QA. Moreover, ArchRAG is very token-efficient, saving up to 250 times the token usage compared to GraphRAG (2024).
@@
  • We present a novel graph-based RAG approach by using ACs that are organized hierarchically and detected by an LLM-based hierarchical clustering method.
@@
  • To index ACs, we propose a novel hierarchical index structure called C-HNSW and also develop an efficient online retrieval method.
@@
  • Extensive experiments show that ArchRAG is both highly effective and efficient, and achieves state-of-the-art performance on both abstract and specific QA tasks.
@@
Figure 2: ArchRAG consists of two phases: offline indexing and online retrieval. For the online retrieval phase, we show an example of using ArchRAG to answer a question in the HotpotQA dataset.
Related Work
$\bullet$ RAG approaches. RAG has been proven to excel in many tasks, including open-ended question answering (2024; 2023), programming context (2024b; 2023), SQL rewrite (2025; 2024), and data cleaning (2024; 2022; 2024). The naive RAG technique relies on retrieving query-relevant contexts from external knowledge bases to mitigate the “hallucination” of LLMs. Recently, many RAG approaches (2025; 2024; 2024; 2024; 2024c; 2024; 2024; 2024) have adopted graph structures to organize the information and relationships within documents, leading to improved performance. For more details, please refer to the recent survey of graph-based RAG methods (2024).
$\bullet$ LLM for graph mining. Recent advances in LLMs have offered opportunities to leverage LLMs in graph mining. These include using LLMs for KG construction (2024), addressing complex graph mining tasks (2024a; 2024; 2024; 2024b), and employing KG to enhance the LLM reasoning (2023; 2024c; 2023; 2023; 2024; 2023; 2025). For instance, RoG (2023) proposes a planning-retrieval-reasoning framework that retrieves reasoning paths from KGs to guide LLMs conducting faithful reasoning. StructGPT (2023) and ToG (2023) treat LLMs as agents that interact with KGs to find reasoning paths leading to the correct answers.
Our Approach ArchRAG
@@
We begin by presenting the overall workflow and design rationale of ArchRAG, followed by detailed descriptions of each component. As illustrated in Figure 2, our proposed ArchRAG consists of two phases. In the offline indexing phase, ArchRAG first constructs a KG from the corpus, then detects ACs by a novel LLM-based hierarchical clustering method, and finally builds the C-HNSW index. During the online retrieval phase, ArchRAG first converts the question into a query vector, then retrieves relevant information from the C-HNSW index, and finally generates answers through an adaptive filtering-based generation process
Our ArchRAG detects ACs by exploiting both links and attributes, and organizes ACs and entities into a novel hierarchical index, C-HNSW, yielding the following advantages: 1) Each AC group densely connected entities with shared themes and a high-quality summary. 2) The hierarchical structure captures multiple levels of abstraction: lower-level entities and communities encode detailed KG knowledge, while higher-level communities provide global context, enabling ArchRAG to address questions at varying granularity. and 3) The C-HNSW index efficiently retrieves relevant information across levels, supporting fast and accurate responses to both abstract and specific questions.
Offline Indexing
KG construction.
@@
ArchRAG builds a KG by prompting the LLM to extract entities and relations from each chunk of the text corpus $D$. Specifically, all text contexts are segmented into chunks based on specified chunk length, enabling the LLM to extract entities and relations from each chunk using in-context learning (2020), thus forming subgraphs. These subgraphs are then merged, with entities and relations that appear repeatedly across multiple subgraphs being consolidated by the LLM to generate a complete description. Finally, we get a KG, denoted by $G(V,E)$, where $V$ and $E$ are sets of vertices and edges, respectively, and each vertex and edge is associated with textual attributes.
LLM-based hierarchical clustering.
@@
We propose an iterative LLM-based hierarchical clustering framework that supports arbitrary graph augmentation (e.g., KNN connections and CODICIL (2013)) and graph clustering algorithms (e.g., weighted Leiden (2019), weighted spectral clustering, and SCAN (2007)). Specifically, we propose to augment the KG by linking entities if their attribute similarities are larger than a threshold, and then associate each pair of linked entities with a weight denoting their attribute similarity value. Next, we generate the ACs using any given graph clustering algorithm. In this way, both node attributes and structural links are jointly considered during community detection.
@@
Algorithm 1 shows the above iterative clustering process. Given a graph augmentation method $Aug$, clustering algorithm $GCluster$ and stopping condition $T$, we perform the following steps in each iteration: (1) augmenting the graph (line 3); (2) computing the edge weights (lines 4-5); (3) clustering the augmented graph (line 6); (4) generating a summary for each community using LLM (line 7); and (5) building a new attributed graph where each node denotes an AC and two nodes are linked if their community members are connected (line 9). We repeat the iterations until the stopping condition $T$ (such as insufficient nodes or reaching the specified level limit) is met. Since each iteration corresponds to one layer, all the ACs $HC$ can be organized into a multi-layer hierarchical tree structure, denoted by $\Delta$, where each community in one layer includes multiple communities in the next layer. Appendix also provides more details.
input : G(V,E){G}(V,E), Aug, GCluster, $T$
@@ 1 $T\leftarrow$ False, ${HC}\leftarrow\emptyset$;
3 ${G}^{\prime}(V,E^{\prime})\leftarrow$ Aug(G(V,E){G}(V,E));
@@
[…]
9 for each $c\in C$ do  generate summary of $c$ by LLM ;